OpenStudy (anonymous):

if "1 > x" is x > 1 or is x < 1 may sound too easy, but this is troubleshooting me a lot.

OpenStudy (amistre64):

if x is less than 1, then x is less than 1 regardless of where you put it

OpenStudy (amistre64):

the symbol is simply defining the relation; it isnt set in stone; if x moves, move the symbol

OpenStudy (amistre64):

think of it as xs little flashlight :) or whatever number/object is using it

OpenStudy (anonymous):

Cannot remember, but some1 told me that in the expression "1>x" if you change the x to the left hand side and the 1 to the left hand side simultaneously, you have to change the direction of the inequality. that is: x<1 ................................................ Today I had to find the domain of the following function: f(x) = Log(Log(x)) I made Log(x) > 0 then I put in into an exponential expression, that is: 10^0 > x 1>x and I flipped the direction of the sign, so I end up with x<0

OpenStudy (amistre64):

i dont think you can create a log(log(x)) function tho

OpenStudy (anonymous):

yeah but you want \[\log(x)>0\] right?

OpenStudy (anonymous):

exactly

OpenStudy (amistre64):

yeah that caveat

OpenStudy (anonymous):

no that is wrong

OpenStudy (amistre64):

there is some trick teachers will play saying find this or that and give a function that makes no sense

OpenStudy (anonymous):

on no \[\ln(\ln(x))\] if a perfectly good function

OpenStudy (anonymous):

I need to find the domain of f(x) = Log(Log(x))

OpenStudy (anonymous):

I know the correct answer is x>0

OpenStudy (anonymous):

an nice example of a function whose limit as x goes to infinity is infinity, but that grows very very VERY slowly

OpenStudy (anonymous):

a no, that is not the right answer

OpenStudy (anonymous):

sorry the correct answer is x>1

OpenStudy (anonymous):

you cannot take the log of a negative number.

OpenStudy (anonymous):

right that is the correct answer

OpenStudy (anonymous):

you have to make sure that \[\log(x)>0\] meaning \[x>1\] done

OpenStudy (amistre64):

for clarity tho http://www.wolframalpha.com/input/?i=domain+log%28log%28x%29%29

OpenStudy (anonymous):

@amistre this is \[\ln(\ln(x))\] and it is graphing real and imaginary parts.

OpenStudy (anonymous):

this is not the "real number" graph.

OpenStudy (amistre64):

wolfram did it, not me :) i cant get it to do what i want it to :)

OpenStudy (anonymous):

the plot is this, right?

OpenStudy (anonymous):

i specified "real" and it gave the domain too!

OpenStudy (anonymous):

probably. it looks like a log

OpenStudy (anonymous):

the picture amistre sent has it, just ignore everything to the left of 1, where it is graphing real and complex parts

OpenStudy (amistre64):

log(x) > 0; when x > 10^0, which is when x>1

OpenStudy (anonymous):

You simplified the log by setting one base in in both sides.

OpenStudy (amistre64):

yeah, i think

OpenStudy (amistre64):

log(1) = 0 ; so everything past it

OpenStudy (anonymous):

I know that way. I want to carry out the the same problem by changing it to the exponential form

OpenStudy (amistre64):

\[log(log(x))=y\] \[10^{log(log(x))}=10^y\] \[log(x)=10^y\] \[10^{log(x)}=10^{10^y}\] \[x =10^{10^y}\] something like that?

OpenStudy (anonymous):

remember that I need to find the domain of the function. THat is , I have to limit the inner log to Log(x) > 0 so I get 10^0 > x. then I get 1 > x. That means that if I want to get to the correct answer I must change x to the left size and the 1 to the right side and no changing the direction of the signg.

OpenStudy (amistre64):

the domain of the function is the range if its inverse :) the range of its inverse is (1,inf)

OpenStudy (amistre64):

\[\lim_{y->-\infty} 10^{10^y}=1\]

OpenStudy (amistre64):

\[log(x)>0\] \[10^{log(x)}>10^0\] \[x>1\] as long as x is greater than 1 your good

OpenStudy (amistre64):

your either a little dyslexic in your process; or your making a simple mistake

OpenStudy (anonymous):

I think the second one haha

OpenStudy (amistre64):

\[log(x)>0\] does not morph into\[10^0>x\]

OpenStudy (anonymous):

that means 1 > x or x > 1, right?

OpenStudy (amistre64):

no; it means that x >1 , which is the same as 1 < x

OpenStudy (amistre64):

1 < x < inf\[\] inf < x < 1

OpenStudy (anonymous):

Im sorry if this is getting boring, but 10^0 > x is the same as 1 > x. Right?

OpenStudy (amistre64):

missed that one lol; ignore the typo

OpenStudy (amistre64):

where are you getting the 10^0 > x from? your inventing that from some nonmath standard and trying to fit it into the problem.

OpenStudy (anonymous):

No. I want to find the domain of the function f(x) = Log(Log(x)). Logs can only be possitive, so I made Log(x)>0.

OpenStudy (amistre64):

good; log(x) > 0 , the next step is?

OpenStudy (anonymous):

put it into an exponetial form. that is 10 ^0 > x, right?

OpenStudy (anonymous):

exponentiate!

OpenStudy (anonymous):

but you have the inequality backwards!

OpenStudy (amistre64):

\[log(x)>0\] invent some new form of non math procedure and come up with \[10^0 > x\] what steps are you using to invent this abomination?

OpenStudy (anonymous):

\[\log(x)>0\] \[10^{\log(x)}>10^0\] \[x>1\]

OpenStudy (anonymous):

i am using "exponentiate"

OpenStudy (amistre64):

use the rules that math provides, and you come to a consistent answer everytime; invent your own rules and you will end up who knows where :)

OpenStudy (anonymous):

the function \[g(x)=10^x\] s a strictly increasing function so if \[a>b\] if follow that \[10^a>10^b\]

OpenStudy (anonymous):

That is what is throublering me. In the equation log(x) = 0, you can do the following: 10^0 = x; can I do the same with inequalities?

OpenStudy (amistre64):

you can as long as you see that you flipped at all around; along with the equals sign

OpenStudy (amistre64):

that sign is as mobile as the others ... it just doesnt make that much of an impact visually

OpenStudy (amistre64):

log(x) = 0\[\] log(x) > 0\[\] 0 = log(x) ; did you see the equal sign flip around too?\[\] 0 < log(x)\[\] 10^0 = x\[\] 10^0 < x\[\]

OpenStudy (amistre64):

make it a general rule; when your flipping things aroung; flip your signs\[\] when you multiply or divide by a "-"; flip your signs .... all of them

OpenStudy (amistre64):

dont try to remember separate rules for separate things when 1 rule does it all

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

yeah finally :) Normally I understand maths quicky, but this time ups ... Thank you very much amistre64 and Satellite

OpenStudy (amistre64):

would you think this is good form?\[\] x < 3\[\] x = 3\[\] 3 = x\[\] 3 < x\[\] ??

OpenStudy (anonymous):

No. I think 3 > x. You have to flip the sign, right?

OpenStudy (amistre64):

you flipped everything around; including the sign ... no matter what the sign looks like, =, >, < , any of them

OpenStudy (anonymous):

so my answer was right?

OpenStudy (amistre64):

yes, your answer was right :)\[\] what property do we have that says:\[\] x = 3 therefore 3=x?\[\] \[x=3\] \[x-3=3-3\] \[x-x-3=0-x\] \[-3=-x\] \[-3(-1)=-x(-1)\text{ ; right here we have to flip the sign right?}\] \[3=x\]

OpenStudy (amistre64):

try it with your > and < signs and see if you need to :)

OpenStudy (anonymous):

how can I pass to the other line? I give enter but everything is shown in one single line

OpenStudy (amistre64):

force a carriage return by putting in an empty equation:\[\] \ then [ then \ then ] \[\] at the end of where you want it to break

OpenStudy (anonymous):

h= 10 +t \ then [ then \ then ]

OpenStudy (anonymous):

???

OpenStudy (amistre64):

take out the "thens" of course ..

OpenStudy (amistre64):

if i try to just type it in to show you it doest\[\]show up ...

OpenStudy (anonymous):

h= 10 +t \ [ \ ]

OpenStudy (amistre64):

so i have to break it apart so it doesnt get feed to the parser as an equation; now remove all the spaces

OpenStudy (anonymous):

ok I think this problem will be solved tomorrow or the day after. So im gonna use mathematica worlfram for the moment. Wait and Im gonna try x=3 with the '>' ;)

OpenStudy (amistre64):

you do that with the >; and it should convince you that your "shorcut" is really just plowing over the real math that is usually done behind the scenes :)

OpenStudy (anonymous):

is that right?